OpenStudy (anonymous):

find an equation for the tangent line to the graph of f at the point P (pi/4, f(pi/4)) when f(x)= 6 tan(x)

OpenStudy (anonymous):

do you know the derivative of \(6\tan(x)\) ?

OpenStudy (anonymous):

6sec^2(x) i think

OpenStudy (anonymous):

Do you know how to find the slope of the tangent line?

OpenStudy (anonymous):

no the teacher is whack and lazy and is being investigated by the university math department

OpenStudy (anonymous):

yes that is right replace \(x\) by \(\frac{\pi}{4}\) and you are done the derivative is the formula for the slope

OpenStudy (anonymous):

the slope of the tangent line is the derivative

OpenStudy (anonymous):

that does not answer what is asked

OpenStudy (anonymous):

you must make an equation

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

Things you need to make an equation of the tangent line

OpenStudy (anonymous):

1. slope of the tangent line

OpenStudy (anonymous):

2. the point which the tangent line passes through

OpenStudy (anonymous):

To get the slope, you take the derivative and plug in your value from your point.

OpenStudy (anonymous):

so plug in \[\frac{ \pi }{ 4 }\] into your derivative. That gives you your slope....

OpenStudy (anonymous):

So, what's your slope?

OpenStudy (anonymous):

12 i think

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

You're saying \[\sec ^{2}(\frac{ \pi }{ 4 }) = 12?\]

OpenStudy (anonymous):

where is the six

OpenStudy (anonymous):

right. so 6sec^2(pi/2) = 12?

OpenStudy (anonymous):

i think im not sure

OpenStudy (anonymous):

haha you're right.

OpenStudy (anonymous):

so. you have your slope.

OpenStudy (anonymous):

ok sorry im slow :(

OpenStudy (anonymous):

now your point is (pi/4, f(pi/4))

OpenStudy (anonymous):

solve for your y in your point.

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

well your point is (x,y) so pi/4=x and f(pi/4)=y, they're basically telling you that your y is the answer you get when you plug pi/4 into your original function.

OpenStudy (anonymous):

right now what

OpenStudy (anonymous):

y=f(pi/4)=6tan(pi/4)

OpenStudy (anonymous):

so solve 6tan(pi/4)